a boomerang is thrown with the following displacements: 4.5-m to the east, 7.5-m 30° north of west, and 3.0-m 10° south of west. what is the resultant displacement?


1
Expert's answer
2020-06-01T14:26:55-0400

Solution.

"AB=4.5m;"

"BC=7.5m;"

"CD=3m;"

"\angle B=30^o;"

"\angle DCB=10^o;"



"AC^2=AB^2+BC^2-2ABBCcos\angle B;"

"AC^2=4.5^2+7.5^2-2\sdot4.5\sdot7.5\sdot 0.866=18.045;"

"AC=4.25m;"

"\dfrac{AC}{sin\angle B}=\dfrac{AB}{sin \angle C}\implies sin\angle C=\dfrac{ABsin\angle B}{AC}" ;


"sin\angle C=\dfrac{4.5\sdot 0.5}{4.25}=0.529;"

"\angle C=32^o;"

"\angle ACD=22^o;"

"AD^2=AC^2+CD^2-2ACCDcos\angle ACD;"

"AD^2=4.25^2+3^2-2\sdot4.25\sdot3\sdot0.9272=3.4189;"

"AD=1.85m;"

"\dfrac{AD}{sin\angle ACD}=\dfrac{CD}{sin \angle CAD}\implies sin\angle CAD=\dfrac{CDsin\angle ACD}{AD};"


"sin\angle CAD=\dfrac{3\sdot 0.3746}{1.85}=0.6074;"

"\angle CAD=37^o;"

"\angle BAD=118^o-37^o=81^o;"

Answer: "AD=1.85m, \angle BAD=81^o;" ,